# Jizhou Liu Booth School of Business

### v1

@@ B.2.1 Sampling Distribution
$\sup_{t\in\mathbb{R}}\left|\Phi\left(\frac{t-\tau}{\sqrt{v_{I}(w^{S})}}\right)-\Phi\left(\frac{t-\tau(w^{S})}{\operatorname{Var}\{T^{B}(\mathbf{W}\mid\mathbf{Y},\mathcal{C})\}^{\frac{1}{2}}}\right)\right|$
$\leq\eta+\frac{|\tau-\tau(w^{S})|+2|\sqrt{v_{I}(w^{S})}-\operatorname{Var}\{T^{B}(\mathbf{W}\mid\mathbf{Y},\mathcal{C})\}^{\frac{1}{2}}|\sqrt{\log(e/\eta)}}{\operatorname{Var}\{T^{B}(\mathbf{W}\mid\mathbf{Y},\mathcal{C})\}^{\frac{1}{2}}}~{}.$
Note that, under $w^{S}\in\mathcal{E}_{1,\eta}$,
$\left|\sqrt{v_{I}(w^{S})}-\operatorname{Var}\{T^{B}(\mathbf{W}\mid\mathbf{Y},\mathcal{C})\}^{\frac{1}{2}}\right|\leq\sqrt{\left|v_{I}(w^{S})-\operatorname{Var}\{T^{B}(\mathbf{W}\mid\mathbf{Y},\mathcal{C})\}\right|}\leq\sqrt{\Delta_{1,\eta}}~{}.$
Therefore, under $w^{S}\in\mathcal{E}_{1,\eta}\bigcap\mathcal{E}_{2,\eta}$ we have
$\sup_{t\in\mathbb{R}}\left|\Phi\left(\frac{t-\tau(w^{S})}{\sqrt{v_{I}(w^{S})}}\right)-\Phi\left(\frac{t-\tau}{\sqrt{v_{I}(w^{S})}}\right)\right|\leq 2\eta+\frac{\Delta_{2,\eta}+4\sqrt{\Delta_{1,\eta}\log(e/\eta)}}{\operatorname{Var}\{T^{B}(\mathbf{W}\mid\mathbf{Y},\mathcal{C})\}^{\frac{1}{2}}}~{}.$
Denote by $\Delta^{\prime}(\eta)$ the summation of the two upper bounds:
$\Delta^{\prime}(\eta)=2\eta+\frac{C\left(\Delta_{I}C_{2}+\sqrt{\Delta_{1,\eta}}\right)}{\operatorname{Var}\{T^{B}(\mathbf{W}\mid\mathbf{Y},\mathcal{C})\}^{\frac{1}{2}}}+\frac{\Delta_{2,\eta}+4\sqrt{\Delta_{1,\eta}\log(e/\eta)}}{\operatorname{Var}\{T^{B}(\mathbf{W}\mid\mathbf{Y},\mathcal{C})\}^{\frac{1}{2}}}~{}.$
Note that $\Delta^{\prime}(\eta)$ is invariant to the seller side treatment allocation $w^{S}$. Next, consider
$\mathbb{E}\sup_{t\in\mathbb{R}}\left|P\left\{\frac{T^{B}(\mathbf{W}\mid\mathbf{Y},\mathcal{C})-\tau}{\sqrt{v_{I}(w^{S})}}\leq t\mid w^{S}\right\}-\Phi\left(t\right)\right|$ $\leq\mathbb{E}\left[1\{w^{S}\in\mathcal{E}_{1,\eta}\cap\mathcal{E}_{2,\eta}\}\Delta^{\prime}(\eta)+1\{w^{S}\not\in\mathcal{E}_{1,\eta}\cap\mathcal{E}_{2,\eta}\}\right]$
$\leq\Delta^{\prime}(\eta)+[1-P(w^{S}\in\mathcal{E}_{1,\eta}\cap\mathcal{E}_{2,\eta})]~{}.$
Following Corollary B.21 by Masoero et al. (2024), there exists a universal constant $C>0$ such that
$\mathbb{E}\sup_{t\in\mathbb{R}}\left|P\left\{\frac{T^{B}(\mathbf{W}\mid\mathbf{Y},\mathcal{C})-\tau}{\sqrt{v_{I}(w^{S})}}\leq t\mid w^{S}\right\}-\Phi\left(t\right)\right|\leq 2\eta+\Delta^{\prime}(\eta)\leq C\Delta\log\left(\frac{C}{\Delta}\right)~{},$ (12)
where
$\Delta:=\frac{C_{1}C_{2}(I^{-1}+J^{-1})}{\operatorname{Var}\{T^{B}(\mathbf{W}\mid\mathbf{Y},\mathcal{C})\}^{\frac{1}{2}}}~{}.$
Under $H_{0}^{wb}:\tau=0$, we have
$\sup_{t\in\mathbb{R}}\left|P\left\{\frac{T^{B}(\mathbf{W}\mid\mathbf{Y},\mathcal{C})}{\sqrt{V^{B}(\mathbf{W}\mid\mathbf{Y},\mathcal{C})}}\sqrt{\frac{v_{I}(w^{S})+S_{\tau}^{2}(w^{S})/I}{v_{I}(w^{S})}}\leq t\right\}-\Phi\left(t\right)\right|$
$=\sup_{t\in\mathbb{R}}\left|\mathbb{E}\left[P\left\{\frac{T^{B}(\mathbf{W}\mid\mathbf{Y},\mathcal{C})}{\sqrt{V^{B}(\mathbf{W}\mid\mathbf{Y},\mathcal{C})}}\sqrt{\frac{v_{I}(w^{S})+S_{\tau}^{2}(w^{S})/I}{v_{I}(w^{S})}}\leq t\mid w^{S}\right\}-\Phi\left(t\right)\right]\right|$
$\leq\mathbb{E}\sup_{t\in\mathbb{R}}\left|P\left\{\frac{T^{B}(\mathbf{W}\mid\mathbf{Y},\mathcal{C})}{\sqrt{v_{I}(w^{S})}}\leq t\mid w^{S}\right\}-\Phi\left(t\right)\right|+\mathbb{E}\sup_{t\in\mathbb{R}}\left|\Phi\left(t\frac{\sqrt{v_{I}(w^{S})+S_{\tau}^{2}(w^{S})/I}}{\sqrt{V^{B}(\mathbf{W}\mid\mathbf{Y},\mathcal{C})}}\right)-\Phi(t)\right|$
$\leq C\Delta\log\left(\frac{C}{\Delta}\right)+\mathbb{E}\sup_{t\in\mathbb{R}}\left|\Phi\left(t\frac{\sqrt{v_{I}(w^{S})+S_{\tau}^{2}(w^{S})/I}}{\sqrt{V^{B}(\mathbf{W}\mid\mathbf{Y},\mathcal{C})}}\right)-\Phi(t)\right|~{}.$
The first inequality follows by Jensen’s inequality and (10). The second inequality follows (12). Then, by dominated convergence theorem and (11), we have
$\lim_{I,J\rightarrow\infty}\mathbb{E}\sup_{t\in\mathbb{R}}\left|\Phi\left(t\frac{\sqrt{v_{I}(w^{S})+S_{\tau}^{2}(w^{S})/I}}{\sqrt{V^{B}(\mathbf{W}\mid\mathbf{Y},\mathcal{C})}}\right)-\Phi(t)\right|=0~{}.$
To interpret the limit of $\Delta$, we follow Masoero et al. (2024) by noting that $\operatorname{Var}\{T^{B}(\mathbf{W}\mid\mathbf{Y},\mathcal{C})\}$ will be of order $(I^{-1}+J^{-1})$. Therefore, we have
$\lim_{I,J\rightarrow\infty}C\Delta\log\left(\frac{C}{\Delta}\right)=0~{}.$
Finally, since $(v_{I}(w^{S})+S_{\tau}^{2}(w^{S})/I)/v_{I}(w^{S})\geq 1$, $T^{WB}(\mathbf{W}\mid\mathbf{Y},\mathcal{C})$ is stochastically dominated by $N(0,1)$.
B.2.2 Randomization Distribution
Next, we show that the randomization distribution of $T^{WB}(\mathbf{W}\mid\mathbf{Y},\mathcal{C})$ converges to $N(0,1)$ in distribution. We denote the randomization distribution of the test statistic and variance estimator with $T^{B,\pi}(\mathbf{W}\mid\mathbf{Y},\mathcal{C})$ and $V^{B,\pi}(\mathbf{W}\mid\mathbf{Y},\mathcal{C})$. Our conditional randomization test takes $\{\bar{Y}_{i}^{B}\}_{i=1}^{I}$ as the fixed input for permuting the treatment vector, and thereby induces a sequence of finite populations $\{\bar{Y}_{i}^{B},\bar{Y}_{i}^{B}\}_{i=1}^{I}$ as the pseudo potential outcomes under both treatment and control. Therefore, the proof should follow exactly Appendix B.2.1. Given a fixed $w^{B}$, with a slight abuse of notation, we should have the same Berry-Esseen type result from Lemma B.5 of Masoero et al. (2024), under $w^{S}\in\mathcal{E}_{1,\eta}$,
$\sup_{t\in\mathbb{R}}\left|P\left\{\frac{T^{B,\pi}(\mathbf{W}\mid\mathbf{Y},\mathcal{C})-\tau(w^{S})}{\sqrt{v_{I}(w^{S})}}\leq t\mid w^{S}\right\}-\Phi\left(t\right)\right|\leq\frac{C\left(\Delta_{I}C_{2}+\sqrt{\Delta_{1,\eta}}\right)}{\operatorname{Var}\{T^{B}(\mathbf{W}\mid\mathbf{Y},\mathcal{C})\}^{\frac{1}{2}}},$
where $\tau(w^{S})=0$ and $v_{I}(w^{S})=\frac{I}{I_{1}I_{0}}S^{2}(w^{S})$ with $S^{2}(w^{S})=\frac{1}{I-1}\sum_{i=1}^{N}(\bar{Y}_{i}^{B}-\bar{Y}^{B})^{2}$ and $\bar{Y}^{B}=\frac{1}{I-1}\sum_{i=1}^{N}\bar{Y}_{i}^{B}$. Note that, under $w^{S}\in\mathcal{E}_{1,\eta}$,
$\sup_{t\in\mathbb{R}}\left|P\left\{\frac{T^{B,\pi}(\mathbf{W}\mid\mathbf{Y},\mathcal{C})}{\sqrt{V^{B,\pi}(\mathbf{W}\mid\mathbf{Y},\mathcal{C})}}\leq t\mid w^{S}\right\}-\Phi\left(t\right)\right|$
$\leq\sup_{t\in\mathbb{R}}\left|P\left\{\frac{T^{B}(\mathbf{W}\mid\mathbf{Y},\mathcal{C})}{\sqrt{v_{I}(w^{S})}}\leq t\mid w^{S}\right\}-\Phi\left(t\right)\right|+\sup_{t\in\mathbb{R}}\left|\Phi\left(t\sqrt{\frac{v_{I}(w^{S})}{V^{B,\pi}(\mathbf{W}\mid\mathbf{Y},\mathcal{C})}}\right)-\Phi(t)\right|~{}.$
The results in Lemma B.1 can be adapted to obtain $V^{B,\pi}(\mathbf{W}\mid\mathbf{Y},\mathcal{C})\xrightarrow{a.s.}v_{I}(w^{S})$ as $I,J\rightarrow\infty$. Then, the rest of the proof should be equivalent to Appendix B.2.1 to finally obtain
$\lim_{I,J\rightarrow\infty}\sup_{t\in\mathbb{R}}\left|P\left\{\frac{T^{B,\pi}(\mathbf{W}\mid\mathbf{Y},\mathcal{C})}{\sqrt{V^{B,\pi}(\mathbf{W}\mid\mathbf{Y},\mathcal{C})}}\leq t\right\}-\Phi\left(t\right)\right|=0~{}.$
B.2.3 Discussions on Total Effects
The proof of validity for total effect tests under the weak null follows the same strategy; thus, we only present the key step. Without loss of generality, we consider a $1$-block conditioning event consisting of diagonal elements from the set of focal units $\{(i,i):w_{i}^{B}=w_{i}^{S}\}$ from a square matrix, i.e., $I=J=n$. Then, the conditional randomization space is defined as $\mathcal{W}=\{\mathbf{W}:w^{B}=w^{S}\}$. The key step in the proof requires showing that
$\lim_{n\rightarrow\infty}\sup_{t\in\mathbb{R}}\left|P\left\{\frac{T^{B}(\mathbf{W}\mid\mathbf{Y},\mathcal{C})-\tau}{\sqrt{v_{I}(\mathcal{W})}}\leq t\right\}-\Phi\left(t\right)\right|=0~{},$ (13)
where $v_{I}(\mathcal{W})$ is defined similarly to $V_{I}(w^{S})$. Consider the conditional probability given as follows:
$\sup_{t\in\mathbb{R}}\left|P\left\{\frac{T^{B}(\mathbf{W}\mid\mathbf{Y},\mathcal{C})-\tau}{\sqrt{v_{I}(\mathcal{W})}}\leq t\mid\mathcal{W}\right\}-\Phi\left(t\right)\right|$
$\leq\sup_{t\in\mathbb{R}}\left|P\left\{\frac{T^{B}(\mathbf{W}\mid\mathbf{Y},\mathcal{C})-\tau(\mathcal{W})}{\sqrt{v_{I}(\mathcal{W})}}\leq t\mid\mathcal{W}\right\}-\Phi\left(t\right)\right|+\sup_{t\in\mathbb{R}}\left|\Phi\left(t-\frac{\tau(\mathcal{W})-\tau}{\sqrt{v_{I}(\mathcal{W})}}\right)-\Phi\left(t\right)\right|~{}.$
By Theorem 1 of Li and Ding (2017), as $n\rightarrow\infty$, the first term converges to 0 under any $\mathcal{W}$. For the second term, by the mean value theorem, there exists $c$ such that
$\sup_{t\in\mathbb{R}}\left|\Phi\left(t-\frac{\tau(\mathcal{W})-\tau}{\sqrt{v_{I}(\mathcal{W})}}\right)-\Phi\left(t\right)\right|\leq\phi(c)\left|\frac{\tau(\mathcal{W})-\tau}{\sqrt{v_{I}(\mathcal{W})}}\right|\leq\frac{1}{\sqrt{2\pi}}\left|\frac{\tau(\mathcal{W})-\tau}{\sqrt{v_{I}(\mathcal{W})}}\right|~{}.$
[…]
B.2.4 Auxiliary Lemmas
Lemma B.1.
Suppose that assumptions in Theorem 4.1 hold, conditional on seller side randomization $w^{S}$, as $I,J\rightarrow\infty$, $V^{B}(\mathbf{W}\mid\mathbf{Y},\mathcal{C})I\xrightarrow{a.s.}v_{I}(w^{S})I+S_{\tau}^{2}(w^{S})$.
@@ B.2.4 Auxiliary Lemmas
Our goal is to show that
$\frac{I}{I_{1}}\frac{1}{I_{1}-1}\sum_{i=1}^{I}w_{i}^{B}\left(\bar{Y}_{i}^{B}-\hat{Y}(1)\right)^{2}+\frac{I}{I_{0}}\frac{1}{I_{0}-1}\sum_{i=1}^{I}(1-w_{i}^{B})\left(\bar{Y}_{i}^{B}-\hat{Y}(0)\right)^{2}\xrightarrow{a.s.}\frac{I}{I_{1}}S_{1}^{2}(w^{S})+\frac{I}{I_{0}}S_{0}^{2}(w^{S})~{}.$
It is sufficient to show that
$\frac{1}{I_{1}-1}\sum_{i=1}^{I}w_{i}^{B}\left(\bar{Y}_{i}^{B}(1)-\hat{Y}(1)\right)^{2}\xrightarrow{a.s.}S_{1}^{2}(w^{S})~{}.$
First, define $\tilde{S}_{z}^{2}(w^{S})=(I_{1}-1)^{-1}\sum_{i=1}^{I}w_{i}^{B}(\bar{Y}_{i}^{B}(z)-\bar{Y}_{\cdot}^{B}(z))^{2}$. Note that we have
$S_{z}^{2}(w^{S})=\frac{1}{I-1}\sum_{i=1}^{I}(\bar{Y}_{i}^{B}(z)-\bar{Y}_{\cdot}^{B}(z))^{2}\leq\frac{2}{I-1}\sum_{i=1}^{I}\bar{Y}_{i}^{B}(z)^{2}+\bar{Y}_{\cdot}^{B}(z)^{2}\leq 4C_{2}^{2}$ (14)
and thus
$\frac{1}{I}\sum_{i=1}^{I}(\bar{Y}_{i}^{B}(z)-\bar{Y}_{\cdot}^{B}(z))^{4}\leq\frac{1}{I}\sum_{i=1}^{I}(\bar{Y}_{i}^{B}(z)^{2}+\bar{Y}_{\cdot}^{B}(z)^{2})^{2}\leq 4C_{2}^{4}~{}.$
By following the proof of Lemma A3 in Wu and Ding (2021), as $I,J\rightarrow\infty$,

### v2

@@ B.2.3 CLT for Neyman-style Studentized Statistic
$\sigma_{A}^{2}:=\lambda\cdot\frac{\pi_{2}}{1-\pi_{2}}\cdot\frac{S_{\mu^{\Delta}}^{2}}{D},\qquad\sigma_{B}^{2}:=1-\frac{S_{\delta}^{2}}{D}.$
Then under Assumptions B.1B.4 and the null hypothesis $H_{0}^{wb,1}:\tau=0$,
$A_{N}\Rightarrow N(0,\sigma_{A}^{2}),\qquad B_{N}\Rightarrow N(0,\sigma_{B}^{2}),\qquad(A_{N},B_{N})\Rightarrow(A,B)\ \text{with }A\perp B,$
and therefore
$T^{WB}(w^{B},w^{S})\Rightarrow N\!\left(0,\ \sigma_{A}^{2}+\sigma_{B}^{2}\right).$
@@
Step 0: Two-stage structure and $\sigma$-fields. Let $\mathcal{G}_{N}:=\sigma(w^{S})$. The seller assignment $w^{S}$ is drawn first; equivalently, $S_{0}(w^{S})$ is an SRSWOR subset of $[J]$ of size $J_{0}$. Conditional on $w^{S}$ (i.e., conditional on $\mathcal{G}_{N}$), the buyer assignment $w^{B}$ is a two-arm complete randomization with fixed margins $(I_{1},I_{0})$ and is independent of $w^{S}$.
Step 1: Unconditional LLN for the denominator $V^{B}(w^{B},w^{S})$.
1(a) SRSWOR mean properties for seller focalization. Fix a buyer $i$ and $z\in\{0,1\}$ and consider the finite population $\{Y_{ij}(z,0)\}_{j=1}^{J}$. Because $S_{0}(w^{S})=\{j:w_{j}^{S}=0\}$ is an SRSWOR subset of size $J_{0}$, the focal mean
$\bar{Y}_{i}^{B}(z;w^{S})=\frac{1}{J_{0}}\sum_{j:w_{j}^{S}=0}Y_{ij}(z,0)$
is an SRSWOR mean. Therefore
$\operatorname*{E}\big[\bar{Y}_{i}^{B}(z;w^{S})\big]=\mu_{i}(z),$
and the SRSWOR variance formula gives
$\operatorname*{Var}\big(\bar{Y}_{i}^{B}(z;w^{S})\big)=\Big(1-\frac{J_{0}}{J}\Big)\frac{S_{i,z}^{2}}{J_{0}},$ (17)
where $S_{i,z}^{2}$ is the finite-population variance over sellers defined in Assumption B.2. Since the SRSWOR mean is unbiased,
$\operatorname*{E}\!\left[\big(\bar{Y}_{i}^{B}(z;w^{S})-\mu_{i}(z)\big)^{2}\right]=\operatorname*{Var}\big(\bar{Y}_{i}^{B}(z;w^{S})\big).$
Combining with (17),
$\operatorname*{E}\!\left[\big(\bar{Y}_{i}^{B}(z;w^{S})-\mu_{i}(z)\big)^{2}\right]=\Big(1-\frac{J_{0}}{J}\Big)\frac{S_{i,z}^{2}}{J_{0}}.$ (18)
1(b) Average $L^{2}$ seller-averaging error vanishes. Average (18) over buyers $i=1,\dots,I$:
$\frac{1}{I}\sum_{i=1}^{I}\operatorname*{E}\!\left[\big(\bar{Y}_{i}^{B}(z;w^{S})-\mu_{i}(z)\big)^{2}\right]$ $=\Big(1-\frac{J_{0}}{J}\Big)\frac{1}{J_{0}}\cdot\frac{1}{I}\sum_{i=1}^{I}S_{i,z}^{2}.$ (19)
By Assumption B.2, $\sup_{N}I^{-1}\sum_{i}S_{i,z}^{2}<\infty$, and $J_{0}\asymp J$, so the right-hand side is $O(J^{-1})\to 0$. Define the nonnegative random variable
$\Delta_{z,N}:=\frac{1}{I}\sum_{i=1}^{I}\big(\bar{Y}_{i}^{B}(z;w^{S})-\mu_{i}(z)\big)^{2}.$
Then $\operatorname*{E}[\Delta_{z,N}]\to 0$ by (19). By Markov’s inequality, for any $\epsilon>0$,
$\Pr(\Delta_{z,N}>\epsilon)\leq\frac{\operatorname*{E}[\Delta_{z,N}]}{\epsilon}\to 0,$
hence
$\Delta_{z,N}\xrightarrow{p}0\qquad(z\in\{0,1\}).$ (20)
1(c) Stability of buyer-level means. By Cauchy–Schwarz,
$\left|\frac{1}{I}\sum_{i=1}^{I}\big(\bar{Y}_{i}^{B}(z;w^{S})-\mu_{i}(z)\big)\right|^{2}\leq\frac{1}{I}\sum_{i=1}^{I}\big(\bar{Y}_{i}^{B}(z;w^{S})-\mu_{i}(z)\big)^{2}=\Delta_{z,N}.$
Thus (20) implies
$\bar{Y}_{\cdot}^{B}(z;w^{S})-\bar{\mu}(z)\xrightarrow{p}0.$
1(d) Stability of second moments. Write
$\bar{Y}_{i}^{B}(z;w^{S})^{2}-\mu_{i}(z)^{2}=\big(\bar{Y}_{i}^{B}(z;w^{S})-\mu_{i}(z)\big)\big(\bar{Y}_{i}^{B}(z;w^{S})+\mu_{i}(z)\big).$
Averaging over $i$ and applying Cauchy–Schwarz gives
$\left|\frac{1}{I}\sum_{i=1}^{I}\bar{Y}_{i}^{B}(z;w^{S})^{2}-\frac{1}{I}\sum_{i=1}^{I}\mu_{i}(z)^{2}\right|$ $\leq\Big(\Delta_{z,N}\Big)^{1/2}\left(\frac{1}{I}\sum_{i=1}^{I}\big(\bar{Y}_{i}^{B}(z;w^{S})+\mu_{i}(z)\big)^{2}\right)^{1/2}.$ (21)
The first factor $\Delta_{z,N}^{1/2}\to_{p}0$ by (20). It remains to show the second factor is $O_{p}(1)$.
Using $(x+y)^{2}\leq 2x^{2}+2y^{2}$,
$\frac{1}{I}\sum_{i=1}^{I}(\bar{Y}_{i}^{B}(z;w^{S})+\mu_{i}(z))^{2}\leq\frac{2}{I}\sum_{i=1}^{I}\bar{Y}_{i}^{B}(z;w^{S})^{2}+\frac{2}{I}\sum_{i=1}^{I}\mu_{i}(z)^{2}.$
By Jensen,
$\mu_{i}(z)^{4}=\Big(\frac{1}{J}\sum_{j=1}^{J}Y_{ij}(z,0)\Big)^{4}\leq\frac{1}{J}\sum_{j=1}^{J}Y_{ij}(z,0)^{4},$
and similarly
$\bar{Y}_{i}^{B}(z;w^{S})^{4}=\Big(\frac{1}{J_{0}}\sum_{j:w_{j}^{S}=0}Y_{ij}(z,0)\Big)^{4}\leq\frac{1}{J_{0}}\sum_{j:w_{j}^{S}=0}Y_{ij}(z,0)^{4}.$
Averaging over $i$ and using Assumption B.1 implies
$\sup_{N}\frac{1}{I}\sum_{i=1}^{I}\mu_{i}(z)^{4}<\infty,\qquad\sup_{N}\operatorname*{E}\!\left[\frac{1}{I}\sum_{i=1}^{I}\bar{Y}_{i}^{B}(z;w^{S})^{4}\right]<\infty.$
Hence, by Markov’s inequality (tightness), both $(I^{-1}\sum_{i}\mu_{i}(z)^{2})$ and $(I^{-1}\sum_{i}\bar{Y}_{i}^{B}(z;w^{S})^{2})$ are $O_{p}(1)$, which shows the second factor in (21) is $O_{p}(1)$. Therefore the left-hand side of (21) converges to 0 in probability:
$\frac{1}{I}\sum_{i=1}^{I}\bar{Y}_{i}^{B}(z;w^{S})^{2}-\frac{1}{I}\sum_{i=1}^{I}\mu_{i}(z)^{2}\xrightarrow{p}0.$ (22)
1(e) Variance stability: $S_{z}^{2}(w^{S})\to_{p}S_{z}^{2}$. Use the identity, valid for any $x_{1},\dots,x_{I}$,
$\frac{1}{I-1}\sum_{i=1}^{I}(x_{i}-\bar{x})^{2}=\frac{I}{I-1}\left(\frac{1}{I}\sum_{i=1}^{I}x_{i}^{2}-\bar{x}^{2}\right).$
Apply it to $x_{i}=\bar{Y}_{i}^{B}(z;w^{S})$ and to $x_{i}=\mu_{i}(z)$ to get
$S_{z}^{2}(w^{S})-S_{z}^{2}$ $=\frac{I}{I-1}\Bigg[\left(\frac{1}{I}\sum_{i=1}^{I}\bar{Y}_{i}^{B}(z;w^{S})^{2}-\bar{Y}_{\cdot}^{B}(z;w^{S})^{2}\right)-\left(\frac{1}{I}\sum_{i=1}^{I}\mu_{i}(z)^{2}-\bar{\mu}(z)^{2}\right)\Bigg]$
$=\frac{I}{I-1}\Bigg[\left(\frac{1}{I}\sum_{i=1}^{I}\bar{Y}_{i}^{B}(z;w^{S})^{2}-\frac{1}{I}\sum_{i=1}^{I}\mu_{i}(z)^{2}\right)-\left(\bar{Y}_{\cdot}^{B}(z;w^{S})^{2}-\bar{\mu}(z)^{2}\right)\Bigg].$
The first bracket converges to 0 in probability by (22). For the second bracket,
$\bar{Y}_{\cdot}^{B}(z;w^{S})^{2}-\bar{\mu}(z)^{2}=\big(\bar{Y}_{\cdot}^{B}(z;w^{S})-\bar{\mu}(z)\big)\big(\bar{Y}_{\cdot}^{B}(z;w^{S})+\bar{\mu}(z)\big).$
The first factor $\to_{p}0$ by Step 1(c); the second factor is $O_{p}(1)$ by the same tightness argument as in Step 1(d). Hence $\bar{Y}_{\cdot}^{B}(z;w^{S})^{2}-\bar{\mu}(z)^{2}\to_{p}0$, and since $I/(I-1)\to 1$ we conclude
$S_{z}^{2}(w^{S})\xrightarrow{p}S_{z}^{2}\qquad(z\in\{0,1\}).$ (23)
1(f) Denominator convergence. Define the finite-population limit
$D:=\frac{S_{1}^{2}}{\pi_{1}}+\frac{S_{0}^{2}}{\pi_{0}}.$
Since $p_{z}=I_{z}/I\to\pi_{z}$ and (23) holds,
$I\left(\frac{S_{1}^{2}(w^{S})}{I_{1}}+\frac{S_{0}^{2}(w^{S})}{I_{0}}\right)=\frac{S_{1}^{2}(w^{S})}{p_{1}}+\frac{S_{0}^{2}(w^{S})}{p_{0}}\xrightarrow{p}\frac{S_{1}^{2}}{\pi_{1}}+\frac{S_{0}^{2}}{\pi_{0}}=D.$
Finally, the sample-variance version $V^{B}(w^{B},w^{S})$ differs from $S_{1}^{2}(w^{S})/I_{1}+S_{0}^{2}(w^{S})/I_{0}$ only by replacing finite-population variances with within-arm sample variances; under complete randomization, these are consistent for their finite-population counterparts, and in particular
$I\,V^{B}(w^{B},w^{S})\xrightarrow{p}D.$ (24)
Step 2: Limit law for $A_{N}=\tau(w^{S})/\sqrt{V^{B}}$.
2(a) Rewrite $\tau(w^{S})$ as an SRSWOR mean over sellers. Let $\Delta_{ij}:=Y_{ij}(1,0)-Y_{ij}(0,0)$ and define seller-level averages
$\mu_{j}^{\Delta}:=\frac{1}{I}\sum_{i=1}^{I}\Delta_{ij}.$
Then
$\tau(w^{S})=\frac{1}{IJ_{0}}\sum_{i=1}^{I}\sum_{j=1}^{J}\Delta_{ij}(1-w_{j}^{S})=\frac{1}{J_{0}}\sum_{j:w_{j}^{S}=0}\mu_{j}^{\Delta}.$
Also,
$\frac{1}{J}\sum_{j=1}^{J}\mu_{j}^{\Delta}=\frac{1}{IJ}\sum_{i=1}^{I}\sum_{j=1}^{J}\Delta_{ij}=\tau.$
Under $H_{0}^{wb}:\tau=0$, the seller-population mean of $\{\mu_{j}^{\Delta}\}$ equals 0.
2(b) Apply the seller-side SRSWOR CLT and rescale. By Lemma B.1,
$\sqrt{J}\,\tau(w^{S})=\sqrt{J}\cdot\frac{1}{J_{0}}\sum_{j:w_{j}^{S}=0}\mu_{j}^{\Delta}\Rightarrow N\!\left(0,\ \frac{\pi_{2}}{1-\pi_{2}}\,S_{\mu^{\Delta}}^{2}\right).$
Now write
$A_{N}=\frac{\sqrt{I}\,\tau(w^{S})}{\sqrt{I\,V^{B}}}=\sqrt{\frac{I}{J}}\cdot\frac{\sqrt{J}\,\tau(w^{S})}{\sqrt{I\,V^{B}}}.$
Using $I/J\to\lambda$ and $I\,V^{B}\to_{p}D$ from (24), Slutsky’s theorem yields